OpenStudy (anonymous):

If x ml pure alcohol is added to 400 ml of a 15% solution and after adding its strength becomes 32%, then the value of x is

OpenStudy (anonymous):

first, convert percents to decimals. Then, use the fact that (concentration1)*(volume1)+(concentration2)*(volume2) = (concentrationF)(volume1+volume2) where volume is measured in milliliters and concentration is the decimal equivalent of percent. I got x=100 mL, which logically makes sense because it has a high concentration but a low amount, so its not going to increase the final concentration by an extreme amount This formula is based on the chemistry formula M1V1=M2V2, relating molarity and volume in acid-base titrations

hero (hero):

Amount Equation: x + 400 = y Alcohol Equation: x + .15(400) = .32y Isolate x in both: x = y - 400 x = .32y - .15(400) Set x = x y - 400 = .32y - 60 Solve for y: y - .32y = 400 - 60 .68y = 340 y = 340/.68 y = 500 solving for x using one of the previous equations: x = y - 400 x = 500 - 400 x = 100

hero (hero):

By the way, this was solved using systems of equations